# Point-JEPA: A Joint Embedding Predictive Architecture for Self-Supervised Learning on Point Cloud

### v3

@@ 3.2 Downstream Tasks
In this section, we report the performance of the learned representation on several downstream tasks. Following the previous studies [27, 29, 7, 28], we report the overall accuracy as a percentage. Unless specified otherwise, we report the mean accuracy and standard deviation from 10 independent runs, each with a different seed, in order to account for variability across independent runs.
@@
After pre-training on ShapeNet [6], we evaluate the learned representation via linear probing on ModelNet40 [26]. Specifically, we freeze the learned context encoder and place the SVM classifier on top. To enforce invariance to geometric transformation, we utilize max and mean pooling on the output of the Transformer encoder [17, 28]. Here, we utilize 1024 points for both training and test sets. As shown in Table 1, our method achieves state-of-the-art accuracy, providing $+0.8\%$ performance gain, showing the robustness of the learned representation.
Table 1: Linear Evaluation on ModelNet40 [26]. We compare Point-JEPA to self-supervised learning methods pre-trained with point cloud data created from ShapeNet [6].
* signifies that the result for linear evaluation is not available in the original paper. We cite the results from [29, 30].
@@
End-to-end Fine-Tuning
We also investigate the performance of the learned representation via end-to-end fine-tuning. After pre-training, we utilize the context encoder to extract the max and average pooled outputs. These outputs are then processed by a three-layer MLP for classification tasks. This class-specific head as well as the context encoder is fine-tuned end-to-end on ModelNet40 [26] and ScanObjectNN [23]. ModelNet40 consists of 12311 synthetic 3D objects from 40 distinct categories, while ScanObjectNN contains objects from 15 classes, each containing 2902 unique instances collected by scanning real-world objects. For ModelNet40, we sub-sample 1024 points per object and sample 64 center points with 32 points in each point patch. On the other hand, we utilize all 2048 points for the ScanObjNN dataset and sample 128 center points with 32 nearest neighbors for the grouped points. As shown in Table 2, our method achieves competitive results when compared to other state-of-the-art methods. Especially, in the OBJ-BG variant of the ScanObjNN [23] dataset, which presents a realistic representation of a point cloud that includes both the object and its background, our method achieves an improvement of $+1\%$ over the best-performing method. This shows the learned representation obtained from pre-training with Point-JEPA can easily be transferred to a classification task.
Table 2: End-to-End Classification. Overall accuracy on ModelNet40 [26] and ScanObjNN [23] with end-to-end fine-tuning. We specifically compare our methods to the method utilizing standard Transformer architecture pre-trained on ShapeNet )[6] with only point cloud (no additional modality.
@@
ModelNet40 ScanObjNN
Method +Voting -Voting OBJ-BG
Point-BERT [27] 93.2 92.7 87.4
Point-MAE [17] 93.8 93.2 90.0
Point-M2AE [29] 94.0 93.4 91.2
Point2Vec [28] 94.8 94.7 91.2
PointGPT-S [7] 94.0 91.6
Point-JEPA (Ours) $94.1\scriptstyle\pm 0.1$ $93.8\scriptstyle\pm 0.2$ 92.9 ±$\scriptstyle\pm 0.4$
Table 3: Result of Few-Shot classification on ModelNet40 [26]. 10 independent trials are completed under one setting. We report mean and standard deviation over $10$ trials.
Overall Accuracy
5-way 10-way
Method 10-shot 20-shot 10-shot 20-shot
@@
We conduct few-shot learning experiments on Modelnet40 [26]. Experiments are done in $m$-way, $n$-shot setting as shown in Table 3. Specifically, we randomly sample $n$ instances of $m$ classes for training. We select 20 instances of $m$ support classes for evaluation. Under one setting, we run 10 independent runs under a fixed random seed on 10 different folds of the dataset, and we report the mean and standard deviation of overall accuracy. As shown in Table 3, our method exceeds the performance of the current state-of-the-art in all settings. Our method yields a $+1.1\%$ improvement in the most difficult 10-way 10-shot setting, showing the robustness of the learned representation of Point-JEPA, especially in the low-data regime.
@@ 4 Conclusion
This work introduced Point-JEPA, a joint embedding predictive architecture applied to point cloud objects. In order to efficiently select targets and context blocks even under the invariance property of point cloud data, we introduced a sequencer, which orders the center points and their corresponding tokens by iteratively selecting the next closest center point. This eliminates the necessity of computing spatial proximity between every pair of tokens or embedding vectors when sampling the targets and context. Our method experimentally achieves state-of-the-art performance in several downstream tasks, showing a strong learned representation. Specifically, Point-JEPA excels in few-shot learning as well as linear evaluation, making the method highly useful when there is a large amount of unlabeled data and a limited amount of labeled data. It is also worth noting that Point-JEPA converges much faster during pre-training, offering a more efficient pre-training alternative in the point cloud domain.

### v4

@@ 4.2 Downstream Tasks
In this section, we report the performance of the learned representation on several downstream tasks. Following the previous studies [40, 42, 8, 41], we report the overall accuracy as a percentage. To account for variability across independent runs, we report the mean accuracy and standard deviation from 10 independent runs with different seeds, unless specified otherwise.
@@
After pre-training on ShapeNet [7], we evaluate the learned representation via linear probing on ModelNet40 [39]. Specifically, we freeze the learned context encoder and place the SVM classifier on top. To enforce invariance to geometric transformation, we utilize max and mean pooling on the output of the Transformer encoder [27, 41]. We utilize 1024 points for both training and test sets. As shown in Table 2, our method achieves state-of-the-art accuracy, providing $+0.8\%$ performance gain, showing the robustness of the learned representation.
Table 2: Linear Evaluation on ModelNet40 [39]. We compare Point-JEPA to self-supervised learning methods pre-trained on ShapeNet [7]. * signifies the linear evaluation results as indicated in [42, 43]. ** signifies results with Transformer backbone.
@@
Table 3: Result of Few-Shot classification on ModelNet40 [39]. 10 independent trials are completed under one setting. We report mean and standard deviation over $10$ trials. ** signifies results with Transformer backbone.
@@
Method 5-way 10-way
10-shot 20-shot 10-shot 20-shot
@@
We conduct few-shot learning experiments on Modelnet40 [39] in $m$-way, $n$-shot setting as shown in Table 3. Specifically, we randomly sample $n$ instances of $m$ classes for training and select 20 instances of $m$ support classes for evaluation. For one setting, we run 10 independent runs under a fixed random seed on 10 different folds of dataset and report mean and standard deviation of overall accuracy. As shown in Table 3, our method exceeds the performance of current state-of-the-art in all settings and yields a $+1.1\%$ improvement in the most difficult 10-way 10-shot setting, showing the robustness of the learned representation of Point-JEPA, especially in the low-data regime.
End-to-end Fine-Tuning
We also investigate the performance of the learned representation via end-to-end fine-tuning. After pre-training, we utilize the context encoder to extract the max and average pooled outputs. These outputs are then processed by a three-layer MLP for classification tasks. This class-specific head as well as the context encoder is fine-tuned end-to-end on ModelNet40 [39] and ScanObjectNN [34]. ModelNet40 consists of 12311 synthetic 3D objects from 40 distinct categories, while ScanObjectNN contains objects from 15 classes, each containing 2902 unique instances collected by scanning real-world objects. For ModelNet40, we sub-sample 1024 points per object and sample 64 center points with 32 points in each point patch. On the other hand, we utilize all 2048 points for the ScanObjNN dataset and sample 128 center points with 32 nearest neighbors for the grouped points.
Table 4: Masking Strategies. Multi-block and single-block masking strategies and their effect on the learned representation.
Targets Context OA
Strategy Ratio Freq. Strategy Ratio Modelnet40 Linear
random (0.6, 0.6) 1 rest 92.5
contiguous (0.6, 0.6) 1 rest 92.3
contiguous (0.15, 0.2) 4 contiguous (0.4, 0.75) 93.7
As shown in Table 1, our method achieves competitive results when compared to other state-of-the-art methods. Especially, in the OBJ-BG variant of the ScanObjNN [34] dataset, which presents a realistic representation of a point cloud that includes both the object and its background, our method achieves the overall accuracy marginally lower than PointDiff [46] while an improvement of $+1\%$ over other SSL methods. This shows the learned representation obtained from pre-training with Point-JEPA can easily be transferred to a classification task.
Table 5: Part Segmentation on ShapeNetPart [7]. mIoUC is the mean IoU for all part categories, and mIoUI is the mean IoU for all instances.
Method mIoUC mIoUI
Transformer-OcCo [40] 83.4 85.1
Point-BERT [40] 84.1 85.6
Point-MAE [27] 84.1 86.1
Point-M2AE [42] 84.9 86.5
Point2Vec [41] 84.6 86.3
PointGPT-S [8] 84.1 86.2
Point-JEPA (Ours) $83.9\pm 0.1$ $85.8\pm 0.1$
Part Segmentation
Following previous studies [40, 42, 41, 27, 8], we report the performance of Point-JEPA in part segmentation task. Here, we utilize the ShapeNetPart [7] dataset, consisting of 16881 objects from 16 categories. We utilize the identical architecture employed in Point2Vec [41] for this task. Specifically, we take the embeddings from $4^{\text{th}}$, $8^{\text{th}}$, and $12^{\text{th}}$ Transformer block and take the average of them. Then, we apply mean and average pooling to this averaged output. The max and mean pooled embedding along with a one-hot encoded class label of an object is used as a global feature vector for the object. The original averaged output is also up-sampled using the PointNet++ [29] feature propagation layer to create a feature vector for each point. Then, each feature vector is concatenated with the global feature vector. A shared MLP is utilized on this concatenated vector to predict the segmentation label for the given point. Although the Point-JEPA shows competitive results, as shown in Table 5, its performance is slightly worse than the state-of-the-art methods.
Limitations
Point-JEPA’s comparatively weaker performance in segmentation and its superior learned representation in classification indicate that the proposed approach emphasizes global features over local features. Additionally, the effectiveness of Point-JEPA for processing larger point clouds is uncertain due to redundancy in many areas in data and requires further study.
4.3 Ablation Study
We conducted thorough ablation studies to understand the effect of moving parts of Point-JEPA. We pre-train Point-JEPA on the ShapeNet [7] dataset under various settings and evaluate the learned representation with linear probing on the ModelNet40 [39] dataset.
Masking Strategy.
We investigate the impact of the masking type on the performance. We consider single-block masking and multi-block masking. For single-block masking strategies, we consider random masking and contiguous masking. For random masking, we randomly select the 60% of indices out of all encoded embedding vectors. Similarly, for contiguous masking, embedding vectors that are spatially contiguous are selected. In this setting, all tokens not corresponding to the selected target blocks are used as context (denoted as rest). On the other hand, in the multi-block masking setting, we sample multiple possibly overlapping spatially contiguous embedding vectors as targets, and we remove the corresponding tokens already utilized for targets during context selection. In this setting, we set the ratio range of 0.15 to 0.2 for targets, while we set the ratio range of 0.4 to 0.75 for context. As shown in Table 4, the single-block masking achieves sub-optimal performance regardless of the spatial contiguity of the target embedding. It shows that our method learns stronger representation by utilizing a smaller amount of targets with a larger frequency.
Table 6: Experiment with different token sequencers using the ModelNet40 linear evaluation and their corresponding training times on RTX 5500 GPU for 500 epochs. “min index” refers to the strategy of selecting the first point in the data feeding order as the initial point. In contrast, “min coordinate” refers to the strategy where the point with the minimum coordinate sum is selected as the initial point.
Token Sequencer Modelnet40 Linear Train Time
z-ordering [24] 93.4 8.30h
Hilbert ordering [33] 91.8 10.78h
Greedy (min index) 92.7 7.47h
Greedy (min coordinate) 93.7 7.47h
Token Sequencers
At the heart of employing Point-JEPA for point cloud processing is the ability to translate point cloud data into a sequence of spatially contiguous tokens. As shown in Table 6, we conduct the ablation experiment with different token sequencer strategies including both the $z$-order [24] and Hilbert-order [33] space-filling curves, as well as the proposed greedy sequencer. The greedy sequencer with the minimum coordinate approach exhibits an overall accuracy improvement compared to all the other algorithms. Additionally, the greedy sequencer offers computational efficiency over z-order and Hilbert-order due to its ability to parallelize computations, as mentioned previously. Notably, the greedy sequencer with minimum coordinate sum initial point guarantees that the starting center point in the sequencer lies near the edge of the object. As shown in Table 6, taking the point near the edge of the object helps with learning stronger representation.
Number of Target Blocks.
We also consider the effect of the number of blocks chosen for targets on the performance of the learned representation while we keep the ratio for targets and context fixed. As shown in Table 7, the performance increases as we increase the number of targets. However, the performance decreases as you increase the number of target blocks after a specific frequency. We observe that our method benefits from having a sufficient amount of tokens available for context encoding.
Table 7: Number of Target blocks. We change the number (frequency) of target blocks while keeping the other components fixed.
Targets Context OA
Ratio Freq. Ratio Modelnet40 Linear
(0.15, 0.2) 1 (0.85, 1.0) 93.0
(0.15, 0.2) 2 (0.85, 1.0) 93.5
(0.15, 0.2) 3 (0.4, 0.75) 93.4
(0.15, 0.2) 4 (0.4, 0.75) 93.7
(0.15, 0.2) 5 (0.4, 0.75) 93.4
(0.15, 0.2) 6 (0.4, 0.75) 93.2
Figure 4: Embedding Visualization on ModelNet40[39]. We visualize the context encoder’s learned representation with t-SNE [35].
4.4 Visualization
To qualitatively analyze the learned representation, we reduce the dimension of the learned representation by utilizing t-SNE [35]. We introduce max and mean pooling on the output of the context encoder, similar to the classification setup, and apply t-SNE on the pooled embedding. We visualize the learned representation on ModelNet40 [39] with no fine-tuning on the dataset. Despite being trained on the dataset, our context encoder produces discriminative features as shown in Figure 4, showing the robustness of the learned representation.
@@ 5 Conclusion
This work introduced Point-JEPA, a joint embedding predictive architecture applied to point cloud objects. In order to efficiently select targets and context blocks even under the invariance property of point cloud data, we introduced a sequencer, which orders the center points and their corresponding tokens by iteratively selecting the next closest center point. This eliminates the necessity of computing spatial proximity between every pair of tokens or embedding vectors when sampling the targets and context. Our method experimentally achieves state-of-the-art performance in several downstream tasks, showing a strong learned representation. Specifically, Point-JEPA excels in few-shot learning as well as linear evaluation, making the method highly useful when there is a large amount of unlabeled data and a limited amount of labeled data. It is also worth noting that Point-JEPA converges much faster during pre-training, offering a more efficient pre-training alternative in the point cloud domain. Future work includes extending Point-JEPA for other downstream tasks such as object detection and scene level segmentation, and pre-training on large unlabeled hybrid datasets [8]. Additionally, the ability of JEPA to reduce dependency on large labeled datasets presents potential avenues the generation of temporal predictive embeddings that anticipate point cloud evolution, thereby enhancing tasks like motion prediction, dynamic scene understanding, and anomaly detection.
Supplementary Material
A Further Pre-training Details
Optimization
We utilize AdamW [22] optimizer with cosine learning decay [21]. Starting from learning rate of $10^{-5}$, we increase it to $10^{-3}$ in the first 30 epochs and decay it to $10^{-6}$. The batch size for pretraining is set to 512, and $\beta$ for Smooth L1 loss is set to $2$, similar to Point2Vec [41]. The target encoder and context encoder initially have identical parameters. The context encoder’s parameters are updated via backpropagation, while the target encoders’ parameters are updated using the exponential moving average of the context encoder parameters, that is $\overline{\theta}\xleftarrow{}\tau\overline{\theta}+(1-\tau)\theta$ where $\tau\in[0,1]$ denotes the decay rate. We gradually increase the decay rate of the exponential moving average from 0.995 to 1.0 during pretraining.
Masking and Ordering
To determine the sequence of tokens, we utilize the iterative ordering of associated centre points, as previously mentioned. We chose the starting point in this sequence with the lowest sum of its coordinates. This method allows us to start the sequence from a point on the outer edge of the object rather than from a point within the object’s interior. This consistency in selecting the initial point is experimentally shown to deliver a slightly better learned representation than taking the first available index.
For masking, we define a range of ratios with both upper and lower limits similar to I-JEPA [2]. To start with, we clarify that the term “block” refers to a sequence of tokens and their corresponding encoded embedding vectors that are contiguous. Because of the sequencing process applied before the target and context selection, most contiguous tokens and encoded embedding vectors are also spatially contiguous. For the target, we randomly select 4 blocks of tokens from within the 0.15 to 0.2 range. We then remove the corresponding tokens of encoded embedding vectors that have already been chosen as targets for further selection. Following this, we choose a block of tokens that is within the range of 0.4 to 0.75 out of available tokens that are not concealed. Because some of the tokens are not available for context selection, we note that context block usually consists of multiple sets of tokens that are spatially contiguous. The selection of targets is completed on a per-batch basis, and we track the indices of these targets to ensure that the corresponding tokens of these selected encoded embedding vectors are concealed in the context selection. The context is then selected using the available indices of tokens also on a per-batch basis.
Targets Context OA
Ratio Freq. Ratio Modelnet40 Linear
(0.1, 0.2) 4 (0.85, 1.0) 93.0
(0.15, 0.2) 4 (0.85, 1.0) 93.3
(0.2, 0.25) 4 (0.85, 1.0) 93.2
(0.25, 0.3) 4 (0.85, 1.0) 92.4
(0.3, 0.35) 4 (0.85, 1.0) 90.5
(0.35, 0.4) 4 (0.85, 1.0) 84.6
Table 8: Ratio Range for Target. The ratio of encoded embedding vectors selected for each target.
B Further Ablation
Ratio of Targets.
We change the ratio of the selected embedding vectors for the target selection while keeping the number of target blocks and the ratio of context tokens fixed. As shown in Table 8, the performance increases when you increase the ratio to a certain point. However, beyond this point, further increasing the ratio results in decreased performance. This implies that Point-JEPA does not require a large size for the target blocks and benefits from a sufficient amount of available tokens for context selection.
Targets Context OA
Ratio Freq. Ratio Modelnet40 Linear
(0.15, 0.2) 4 (0.85, 1.0) 93.1
(0.15, 0.2) 4 (0.75, 1.0) 92.8
(0.15, 0.2) 4 (0.65, 1.0) 93.4
(0.15, 0.2) 4 (0.45, 1.0) 93.6
(0.15, 0.2) 4 (0.6, 0.75) 93.4
(0.15, 0.2) 4 (0.5, 0.75) 93.1
(0.15, 0.2) 4 (0.4, 0.75) 93.7
Table 9: Ratio Range for Context. The ratio of tokens selected for context encoding.
(a) Row-normalized confusion matrix on ModelNet40
(b) Column-normalized confusion matrix on ModelNet40
(c) Row-normalized confusion matrix on ScanObjNN
(d) Column-normalized confusion matrix on ScanObjNN
Figure 5: Confusion matrices illustrating model performance on ModelNet40 and another dataset, highlighting class-specific accuracies and challenges with similar categories.
Ratio of Context.
In this study, we change the ratio of tokens selected for context encoding while keeping the number of targets and the ratio range for targets fixed. As shown in Table 9, having a relatively large difference between the lower and upper bound of the ratio can improve performance. In other words, Point-JEPA learns a better representation when the number of selected context tokens varies more between training iterations. Additionally, when the upper bound of the ratio is somewhat constrained, we see increased performance.
Predictor Depth
We also study the effect of the predictor’s depth on the learned representation. To this end, we vary the predictor depth and observe its effect on the linear evaluation accuracy. As shown in Table 10, Point-JEPA benefits from a deeper predictor.
Predictor Depth Modelnet40 Linear (OA)
2 92.5
3 92.8
4 93.2
5 93.4
6 93.7
Table 10: Predictor Depth. Predictor depth and its effect on learned representation.
Class confusion on ModelNet40 and ScanObjNN
To assess our model’s performance on the ModelNet40 [39] and ScanObjNN [34] datasets, we present two types of visualizations for each dataset. The first is a row-normalized confusion matrix, which illustrates the model’s sensitivity, indicating how well the model identifies each actual class. The second is a column-normalized confusion matrix, depicting the model’s specificity, which shows the correctness of predictions for each class assumed by the model. As illustrated in parts (a) and (b) of Fig. 5, the model fine-tuned on ModelNet40 demonstrates high accuracy. At the same time, errors predominantly arise from similar categories within the dataset. For instance, “flower pot” and “plant” are often misclassified, likely due to the presence of flowers in some of the flower pot models in the ModelNet40 dataset. Similarly, parts (c) and (d) of Fig. 5 show the aforementioned confusion matrices. As highlighted in the main paper, our model’s performance on ScanObjectNN dataset has room for enhancement compared to ModelNet40. The confusion matrix reveals some misclassifications, but it is encouraging to see that these errors predominantly occur between closely related classes, such as ‘table’ and ‘desk’ or ‘sofa’ and ‘bed’. This suggests that our model has a solid grasp of the key characteristics of these categories and that further refinement of the classification criteria could lead to significant improvements in overall accuracy.